OpenStudy (anonymous):

Find the sum of all 2 digit numbers N=ab¯¯¯, where a≠0, such that N divides a0b¯¯¯¯¯. Clarification: The notation a0b¯¯¯¯¯ used in the question means that we read it as a 3-digit number. For example, with a=4,b=8, a08¯¯¯¯¯ represents the number 408, and not the number 4×0×8=0.

OpenStudy (mayankdevnani):

(100a+b)/(10a+b)=n where 'n' is integer. Let, b=0 so, 100a/10a=10 here 'a' can take 9 values from 1 to 9 so, we get 10+20+30.....+90=450 'a' not equal to 'b' cause 101a/11a is not an integer. To find more values, we have to do some calculations. 100a+b=10na+nb =>10a(10-n)=b(n-1) =>10a/b=(n-1)/(10-1) when n=6 10a/b=5/4 =>40a=5b =>8a=b =>a/b=1/8 so, a=1 , b=8 (a,b can not take value greater than 9). Get the number 18 in this way, trying with n=7, 9 get the numbers 15, 45 (i tried the values 1 to 9 and got for only n=6,7,9) so the sum is 450+15+18+45=528